# Rayleigh-Bénard convective motion of stratified fluids in the Earth’s troposphere

### v1

@@ II.2 Convective flow in a bounded domain
Comparing this expression with that in the unbounded case [Case III, Eq. (16) for $\gamma>0$], we find that the wave number $k$ is now replaced by $m^{2}+k^{2}_{\perp}$, i.e., the vertical wave number $k_{z}$, which assumes any values in the unbounded domains (See Case III of Sec. II.1), is replaced by the discrete eigenvalue $m$, and estimating the instability growth rate becomes an eigenvalue problem. Similar to Eq. (16), the minimum or critical value of the Rayleigh number above which the convective instability can occur is given by [See also Eq. (41)]
@@
where $m=(2n+1)\pi/H$ with $n=0,\pm 1,\pm 2,...$ and the smallest critical value attained at the vertical wave number $mH=\pm\pi$ (corresponding to $n=0$ or $n=-1$) is approximately $657.5$ [See also Eq. (42)]. It is important to note that, in contrast to the case of unbounded region, the wave number $mH$ does not assume arbitrary values, but discrete eigenvalues corresponding to the integer $n$. The qualitative features of the instability growth rate will remain the same as Fig. 1, and we do not repeat them here for brevity.
@@
This case is similar to Case IV of Sec. II.1 in which the effects of the BV frequency and dissipation are retained, however, in a bounded region. Substitution of Eq. (40) into Eq. (32) gives
@@
where we have considered the plus sign before the square root, and since $m\tan(mz)=-\left[1/\tilde{p}(z)\right]\left[d\tilde{p}(z)/dz\right]=1/L_{\tilde{p}}>0$ for $d\tilde{p}(z)/dz<0$ (Typically, the atmospheric pressure decreases with increasing the altitude and $L_{\tilde{p}}$ denotes the vertical length scale of pressure inhomogeneity) and $N^{2}<0$ (which holds for the instability of stratified fluids [6, 7]), we have $\widetilde{N}^{2}>0$. Thus, $k^{2}_{\perp}>0$ and the right-hand side of Eq. (63) is also positive.
Inserting Eq. (64) in the right-hand side of Eq. (63), we find
@@
Figure 5 shows the profiles of the instability growth rate $\gamma$, given by Eq. (61), for different values of the dimensionless parameters: $\beta_{1}\equiv\beta T_{0}$, $\chi_{1}\equiv\chi/\sqrt{gH^{3}}$, $\omega_{g}^{2}=-N^{2}H/g$, and $L_{\tilde{p}}/H$ (when $d\tilde{p}(z)/dz<0$) as mentioned in the figure caption. For a fixed set of parameter values, the qualitative feature of the growth rate is similar to Fig. 3(b). We note that the growth rate is significantly enhanced by the effects of increasing values of the thermal expansion parameter $\beta$ and the vertical length scale of pressure inhomogeneity $L_{\tilde{p}}$ compared to that of perturbations $H$ (Compare the solid line with dashed and thick dashed lines). Thus, in contrast to the unbounded domain, the background pressure inhomogeneity effect favors the convective instability with higher growth rates as the corresponding scale size tends to increase compared to the vertical scale of perturbations. Furthermore, the thermal expansion results in a higher growth rate with cut-offs at higher wave numbers compared to the unbounded domain. While the effect of the BV frequency is similar to the bounded domain in reducing the growth rate with its increasing values, the dissipation due to the kinematic viscosity and thermal diffusivity does not have any significant effect in increasing or decreasing the maximum growth rate but giving cut-offs at lower wave numbers as the value of $\chi$ increases (analogous to the case of the bounded domain). The effects of the other parameters on the growth rate are not discussed for the reason mentioned in the figure caption, or in the discussion the discussion of Figs. 1 and 3.
@@
Figure 5: The instability growth rate [$\gamma$, given by Eq. (61),is shown (Case II, Sec. II): $\chi\neq 0,~{}\nu\neq 0,~{}N^{2}\neq 0$) for different values of the parameters: $\beta_{1}\equiv\beta T_{0}$, $\chi_{1}\equiv\chi/\sqrt{gH^{3}}$, $\omega_{g}^{2}=-N^{2}H/g$, and $L_{\tilde{p}}$. The solid, dashed, dotted, dash-dotted, and thick dashed lines, respectively, correspond to the values (i) $\beta_{1}=0.25,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.24$, and $k_{z}H=0.3$, (ii) $\beta_{1}=0.5,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.24$, and $L_{\tilde{p}}=5$, (iii) $\beta_{1}=0.25,~{}\chi_{1}=0.06,~{}\omega_{g}^{2}=0.24$, and $L_{\tilde{p}}/H=5$, (iv) $\beta_{1}=0.25,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.41$, and $L_{\tilde{p}}=5$, and (v) $\beta_{1}=0.25,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.24$, and $L_{\tilde{p}}/H=10$. Here, $H$ is the vertical length scale of perturbation and $L_{\tilde{p}}$ the vertical length scale of pressure inhomogeneity. Note that the parameters $\nu$ and $\Gamma$ have the similar effects on the instability growth rate as of $\chi$ and $\beta$ respectively (not shown). The fixed parameter values are the same as for Fig. 1, i.e., $H=800$ m, $g=9.8~{}\rm{m/s}^{2}$, $T_{0}=250$ K, $\Gamma=0.0065$ K/m, and $\nu_{1}\equiv\nu/\sqrt{gH^{3}}=0.03$ together with $m=\pi/H$ at $n=0$. The critical values corresponding to the solid, dashed, dotted lines are the same, i.e., $Ra_{\min}=659.7$, and to the dash-dotted and thick dashed lines, respectively, are $Ra_{\min}=661.1$ and $658.6$. For $m=3\pi/H$ at $n=1$, and other parameter values as for the solid line, we have $Ra_{\min}=5.3\times 10^{4}$.
III Summary and Conclusion
@@
We have studied the Rayleigh-Bénard convective wave motion and instability of stratified fluids under gravity in the troposphere as motivated by the recent work of Kaladze and Misra [7] in which they showed that the effects of temperature-dependent density inhomogeneity due to thermal expansion can significantly modify the Brunt-Väisälä (BV) frequency and lead to the onset of convective instability with negative thermal gradient. However, they restricted the model to an unbounded domain of the troposphere and did not explore or discuss many relevant details. We have revisited the theory of convective motion and instability in more detail by considering both unbounded and bounded tropospheric domains to show that the conditions for instability in these two cases significantly differ. Thus, the previous theory of convective instability in an unbounded domain is generalized and advanced. The critical values of the Raleigh numbers and the expressions for instability growth rates of thermal waves are obtained and analyzed in various parameter regimes of dissipation and thermal expansion associated with the BV frequency in both unbounded and bounded domains. In the latter, we get the necessary boundary conditions and show the quantification of the vertical wave number and the corresponding eigenvalue problem. The main findings can be summarized as follows:
@@
    In the most simple case, when $\chi=\nu=N^{2}=0$, convective waves without any damping or instability can propagate only in the stratosphere when $\Gamma<0$, i.e., with positive thermal gradient. However, in the troposphere with $\Gamma<0$, convective instability can occur with a zero real wave frequency but a finite growth rate. The instability becomes stronger and the growth rate becomes higher (without any cut-off) as the vertical scale $H$ or the wavelengths of perturbations gradually increase or the thermal expansion coefficient $\beta$ tends to increase.
@@
    In the case, when $\chi=\nu=0$ but $N^{2}\neq 0$, the propagation of convective waves with a larger growth rate than the real frequency is possible (i.e., $\gamma>\omega_{r}$). Such growth rates can be associated with the aperiodic fluid motion and higher values of the Rayleigh number where dominant effects of the buoyancy force over dissipation come into the picture. Similar to item (i), the instability growth rate increases with increasing values of $H$ or $\beta$.
@@
    In contrast to the case of item (ii), i.e., $\chi\neq 0,~{}\nu\neq 0$ but $N^{2}=0$, the convective instability with a zero real frequency but a finite growth rate (i.e., $\omega_{r}=0,~{}\gamma>0$) can occur. The existence of zero-frequency waves with $\omega_{r}=\gamma=0$ is also possible at some critical value of the wave number, i.e., they appear as a flat line and may exist everywhere instantly. Furthermore, a minimum, or a critical value of the Rayleigh number, at which the convection starts and above which the instability sets in, is obtained and shown to be proportional to the vertical wave number $k_{z}H$. It is noted that as the vertical length scale of perturbations $H$ deepen, or the value of $\chi$ or $\nu$ is reduced, or that of $\beta$ is increased, the convective wave tends to become more unstable with a higher growth rate.
@@
    In a general situation when $\chi\neq 0,~{}\nu\neq 0$, and $N^{2}\neq 0$, convective wave motion and instability are possible by the influence of the BV frequency with $\omega_{r},~{}\gamma>0$. The critical or minimum value of the Rayleigh number gradually increases with the horizontal wave number $k_{\perp}H$, and as it increases, the growth rate tends to vanish after achieving a maximum value, implying that the larger the minimum value of the Rayleigh number, the less unstable is the wave motion.
@@
    In the case of $\chi\neq 0,~{}\nu\neq 0$ but $N^{2}=0$, we have the possibility of the convective instability with $\gamma>0$ and $\omega_{r}=0$, similar to item (iii) above for an unbounded domain. However, in contrast to (iii), the vertical wave number no longer assumes any values but some discrete eigenvalues. The minimum value of the corresponding Rayleigh number is found to be approximately $658$, which is much higher than that can be predicted in an unbounded domain.
@@
    when $\chi\neq 0,~{}\nu\neq 0$, and $N^{2}\neq 0$, we have also the possibility of the convective instability with $\gamma>0$ and $\omega_{r}=0$. However, in addition to the BV frequency $N^{2}$ and the vertical wave number $m$ (eigenvalue), a new vertical length scale of pressure inhomogeneity $(L_{\tilde{p}})$ appears, which significantly influences the minimum value of the Rayleigh number and the instability growth rate. It is noted that the background pressure inhomogeneity effect favors the convective instability with higher growth rates as the length scale $L_{\tilde{p}}$ tends to increase compared to the vertical scale of perturbations $(H)$.
To conclude, The present investigation is an attempt to successively improve the existing theory of wave motion and instability of atmospheric stratified fluids under gravity. In this way, it advances the previous works [6, 7]. The wave motions and instabilities reported here could help better understand different types of weather systems and their severity in unstable conditions compared to the previous studies. The authors look for further advancement in a medium of rotating fluids [9, 10, 11].
Conflict of Interest
The authors have no conflicts to disclose.
Data availability statement
All data that support the findings of this study are included within the article (and any supplementary files).

### v2

@@ 2.2 Convective flow in a bounded domain
Comparing this expression with that in the unbounded case [Case III, Eq. (16) for $\gamma>0$], we find that the wave number $k$ is being replaced by $m^{2}+k^{2}_{\perp}$, i.e., the vertical wave number $k_{z}$, which assumes any values in the unbounded domains (See Case III of Sec. 2.1), is replaced by the discrete eigenvalue $m$, and estimating the instability growth rate becomes an eigenvalue problem. Similar to Eq. (16), the minimum or critical value of the Rayleigh number above which the convective instability can occur is given by [See also Eq. (41)]
@@
where $m=(2n+1)\pi/H$ with $n=0,\pm 1,\pm 2,...$ and the smallest critical value attained at the vertical wave number $mH=\pm\pi$ (corresponding to $n=0$ or $n=-1$) is approximately $657.5$ [See also Eq. (42)]. It is important to note that, in contrast to the case of the unbounded region, the wave number $mH$ does not assume arbitrary values but discrete eigenvalues corresponding to the integer $n$. The qualitative features of the instability growth rate will remain the same as Fig. 1, and we do not repeat them here for brevity.
@@
This case is similar to Case IV of Sec. 2.1 in which the effects of the BV frequency and dissipation are retained, however, in a bounded region. Substitution of Eq. (40) into Eq. (32) gives
@@
where we have considered the plus sign before the square root, and since $m\tan(mz)=-\left[1/\tilde{p}(z)\right]\left[d\tilde{p}(z)/dz\right]=1/L_{\tilde{p}}>0$ for $d\tilde{p}(z)/dz<0$ (Typically, the atmospheric pressure decreases with increasing the altitude and $L_{\tilde{p}}$ denotes the vertical length scale of pressure inhomogeneity) and $N^{2}<0$ (which holds for the instability of stratified fluids (Kaladze and Misra, 2023, 2024)), we have $\widetilde{N}^{2}>0$. Thus, $k^{2}_{\perp}>0$ and the right-hand side of Eq. (63) is also positive. Inserting Eq. (64) in the right-hand side of Eq. (63), we find
@@
Figure 5 shows the profiles of the instability growth rate $\gamma$, given by Eq. (61), for different values of the dimensionless parameters: $\beta_{1}\equiv\beta T_{0}$, $\chi_{1}\equiv\chi/\sqrt{gH^{3}}$, $\omega_{g}^{2}=-N^{2}H/g$, and $L_{\tilde{p}}/H$ (when $d\tilde{p}(z)/dz<0$) that are relevant for tropospheric fluids (Kaladze and Misra, 2023) as mentioned in the figure caption. For a fixed set of parameter values, the qualitative feature of the growth rate is similar to Fig. 3(b). We note that the growth rate is significantly enhanced by the effects of increasing values of the thermal expansion parameter $\beta$ and the vertical length scale of pressure inhomogeneity $L_{\tilde{p}}$ compared to that of perturbations $H$ (Compare the solid line with dashed and thick dashed lines). Thus, in contrast to the unbounded domain, the background pressure inhomogeneity effect favors the convective instability with higher growth rates as the corresponding scale size tends to increase compared to the vertical scale of perturbations. Furthermore, the thermal expansion gives rise to a higher growth rate with cut-offs at higher wave numbers relative to the unbounded domain. While the effect of the BV frequency is similar to the bounded domain in reducing the growth rate with its increasing values, the dissipation due to the kinematic viscosity and thermal diffusivity does not have any significant effect in increasing or decreasing the maximum growth rate but giving cut-offs at lower wave numbers as the value of $\chi$ increases (analogous to the case of the bounded domain). We do not discuss the effects of the other parameters on the growth rate for the reason either mentioned in the figure caption or the discussion of Figs. 1 and 3.
@@
Figure 5: The instability growth rate [$\gamma$, given by Eq. (61),is shown (Case II, Sec. 2): $\chi\neq 0,~{}\nu\neq 0,~{}N^{2}\neq 0$) for different values of the parameters: $\beta_{1}\equiv\beta T_{0}$, $\chi_{1}\equiv\chi/\sqrt{gH^{3}}$, $\omega_{g}^{2}=-N^{2}H/g$, and $L_{\tilde{p}}$. The solid, dashed, dotted, dash-dotted, and thick dashed lines, respectively, correspond to the values (i) $\beta_{1}=0.25,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.24$, and $k_{z}H=0.3$, (ii) $\beta_{1}=0.5,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.24$, and $L_{\tilde{p}}=5$, (iii) $\beta_{1}=0.25,~{}\chi_{1}=0.06,~{}\omega_{g}^{2}=0.24$, and $L_{\tilde{p}}/H=5$, (iv) $\beta_{1}=0.25,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.41$, and $L_{\tilde{p}}=5$, and (v) $\beta_{1}=0.25,~{}\chi_{1}=0.04,~{}\omega_{g}^{2}=0.24$, and $L_{\tilde{p}}/H=10$. Here, $H$ is the vertical length scale of perturbation and $L_{\tilde{p}}$ the vertical length scale of pressure inhomogeneity. Note that the parameters $\nu$ and $\Gamma$ have the similar effects on the instability growth rate as of $\chi$ and $\beta$ respectively (not shown). The fixed parameter values are the same as for Fig. 1, i.e., $H=800$ m, $g=9.8~{}\rm{m/s}^{2}$, $T_{0}=250$ K, $\Gamma=0.0065$ K/m, and $\nu_{1}\equiv\nu/\sqrt{gH^{3}}=0.03$ together with $m=\pi/H$ at $n=0$. The critical values corresponding to the solid, dashed, dotted lines are the same, i.e., $Ra_{\min}=659.7$, and to the dash-dotted and thick dashed lines, respectively, are $Ra_{\min}=661.1$ and $658.6$. For $m=3\pi/H$ at $n=1$, and other parameter values as for the solid line, we have $Ra_{\min}=5.3\times 10^{4}$.
3 Summary and Conclusion
@@
We have studied the Rayleigh-Bénard convective wave motion and instability of stratified fluids under gravity in the troposphere as motivated by the recent work of Kaladze and Misra (Kaladze and Misra, 2024) in which they showed that the effects of temperature-dependent density inhomogeneity due to thermal expansion can significantly modify the Brunt-Väisälä (BV) frequency and lead to the onset of convective instability with negative thermal gradient. However, they restricted the model to an unbounded domain of the troposphere and did not explore or discuss many relevant details. We have revisited the theory of convective motion and instability in more detail by considering both unbounded and bounded tropospheric domains to show that the conditions for instability in these two cases significantly differ. Thus, the previous theory of convective instability in an unbounded domain is generalized and advanced. The critical values of the Raleigh numbers and the expressions for instability growth rates of thermal waves are obtained and analyzed in various parameter regimes of dissipation and thermal expansion associated with the BV frequency in both unbounded and bounded domains. In the latter, we get the necessary boundary conditions and show the quantification of the vertical wave number and the corresponding eigenvalue problem. The main findings can be summarized as follows:
@@
    In the simple case, when $\chi=\nu=N^{2}=0$, convective waves without any damping or instability can propagate only in the stratosphere when $\Gamma<0$, i.e., with a positive thermal gradient. However, in the troposphere with $\Gamma<0$, convective instability can occur with a zero real wave frequency but a finite growth rate. The instability becomes stronger, and the growth rate becomes higher (without any cut-off) as the vertical scale $H$ or the wavelengths of perturbations gradually increase, or the thermal expansion coefficient $\beta$ tends to increase.
@@
    When $\chi=\nu=0$ but $N^{2}\neq 0$, the propagation of convective waves with a larger growth rate than the real frequency is possible (i.e., $\gamma>\omega_{r}$), and this growth rate can be associated with the aperiodic fluid motion and higher values of the Rayleigh number, where dominant effects of the buoyancy force over dissipation come into the picture. Like item (i), the instability growth rate increases with increasing values of $H$ or $\beta$.
@@
    In contrast to the case of item (ii), i.e., $\chi\neq 0,~{}\nu\neq 0$ but $N^{2}=0$, the convective instability with a zero real frequency but a finite growth rate (i.e., $\omega_{r}=0,~{}\gamma>0$) can occur. The existence of zero-frequency waves with $\omega_{r}=\gamma=0$ is also possible at some critical value of the wave number, i.e., they appear as a flat line and may exist everywhere instantly. Furthermore, a minimum, or a critical value of the Rayleigh number, at which the convection starts and above which the instability sets in, is obtained and shown to be proportional to the vertical wave number $k_{z}H$. We note that as the vertical length scale of perturbations $H$ deepen, or the value of $\chi$ or $\nu$ gets reduced, or that of $\beta$ increases, the convective wave tends to become more unstable with a higher growth rate.
@@
    In a general situation when $\chi\neq 0,~{}\nu\neq 0$, and $N^{2}\neq 0$, convective wave motion and instability are possible by the influence of the BV frequency with $\omega_{r},~{}\gamma>0$. The critical or minimum value of the Rayleigh number gradually increases with the horizontal wave number $k_{\perp}H$, and as its value increases, the growth rate tends to vanish after achieving a maximum value, implying that the larger the minimum value of the Rayleigh number, the less unstable is the wave motion.
@@
    In the case of $\chi\neq 0,~{}\nu\neq 0$ but $N^{2}=0$, we have the possibility of the convective instability with $\gamma>0$ and $\omega_{r}=0$, similar to item (iii) above for an unbounded domain. However, in contrast to (iii), the vertical wave number no longer assumes any values but some discrete eigenvalues. The minimum value of the corresponding Rayleigh number is approximately $658$, which is much higher than predicted in an unbounded domain.
@@
    When $\chi\neq 0,~{}\nu\neq 0$, and $N^{2}\neq 0$, we also have the possibility of convective instability with $\gamma>0$ and $\omega_{r}=0$. However, in addition to the BV frequency $N^{2}$ and the vertical wave number $m$ (eigenvalue), a new vertical length scale of pressure inhomogeneity $(L_{\tilde{p}})$ appears, which significantly influences the minimum value of the Rayleigh number and the instability growth rate. We have noted that the background pressure inhomogeneity effect favors the convective instability with higher growth rates as the length scale $L_{\tilde{p}}$ tends to increase compared to the vertical scale of perturbations $(H)$.
To conclude, The present investigation is an attempt to successively improve the existing theory of wave motion and instability of atmospheric stratified fluids under gravity. In this way, it advances the previous works (Kaladze and Misra, 2023, 2024). The wave motions and instabilities reported here could help better understand different types of weather systems and their severity in unstable conditions compared to the previous studies. The authors look for further advancement in a medium of rotating fluids (Chatterjee and Misra, 2021; Abdikian et al., 2022; Turi and Misra, 2022).
Data availability
Data will be made available on request.